The difference between a grandmother's and granddaughter's age is 44 years. If the grandmother's age is 13 less than 4 times the age of the granddaughter, how old is each now?
Answer by lwsshak3(11628)

The difference between a grandmother's and granddaughter's age is 44 years. If the grandmother's age is 13 less than 4 times the age of the granddaughter, how old is each now?
***
let x=granddaughter's age
x+44=grandmother's age
..
x+44=4x-13
3x=57
x=19
x+44=63
granddaughter's age=19
grandmother's age=63